# It’s TIME: Towards the Next Generation of Time Series Forecasting Benchmarks

### v1

@@ Appendix E Full Results
Coastal T-S/15T MASE (Raw) 1.261 1.576 1.263 1.277 1.299 1.259 1.303 1.272 1.284 1.563 1.322 1.291
CRPS (Raw) 0.072 0.069 0.081 0.070 0.076 0.068 0.069 0.093 0.080 0.075 0.069 0.082
MASE (Norm.) 0.819 1.026 0.821 0.830 0.846 0.818 0.845 0.827 0.834 1.019 0.858 0.839
CRPS (Norm.) 0.666 0.631 0.742 0.641 0.700 0.628 0.634 0.857 0.725 0.692 0.637 0.750
MASE (Rank) 1.667 12.333 3.333 5.333 7 2 8.667 4 6.333 12 9.333 7.333
CRPS (Rank) 6.667 2.333 10 4.667 9.333 1.333 2.667 11.333 7.667 8 4 10
Coastal T-S/20T MASE (Raw) 0.881 0.996 1.271 1.131 1.276 1.116 1.102 1.488 1.268 1.712 2.107 1.074
CRPS (Raw) 0.004 0.004 0.006 0.005 0.006 0.006 0.005 0.006 0.005 0.009 0.011 0.005
MASE (Norm.) 0.547 0.640 0.809 0.715 0.878 0.743 0.673 0.873 0.762 1.116 1.258 0.674
CRPS (Norm.) 0.512 0.601 0.877 0.708 0.995 0.896 0.710 0.860 0.717 1.474 1.456 0.697
MASE (Rank) 1 3.333 8.667 6.333 7.333 5.667 4 9 7.333 11.667 12 4.667
CRPS (Rank) 1 2.667 9.333 6 8 6 5.333 9.667 5.667 12.333 11.667 5
Coastal T-S/H MASE (Raw) 1.618 1.999 1.732 1.637 1.797 1.718 1.602 1.955 1.859 2.106 1.896 1.909
CRPS (Raw) 0.029 0.030 0.033 0.030 0.033 0.031 0.029 0.034 0.035 0.031 0.033 0.038
MASE (Norm.) 0.778 1.001 0.842 0.791 0.902 0.820 0.774 0.925 0.888 1.054 0.920 0.920
CRPS (Norm.) 0.576 0.603 0.648 0.597 0.667 0.617 0.576 0.682 0.679 0.622 0.667 0.739
MASE (Rank) 2.333 9 5.667 3 6 4.333 2 10 7 12.333 8.667 9.333
CRPS (Rank) 2 4.333 7.667 3.667 8 5 1.333 10.667 9 6 9.333 11
SG Weather MASE (Raw) 0.881 0.882 0.899 0.899 1.037 0.889 0.888 0.906 0.913 0.906 0.961 0.932
CRPS (Raw) 0.177 0.176 0.181 0.185 0.215 0.178 0.178 0.200 0.188 0.187 0.201 0.196
MASE (Norm.) 0.739 0.740 0.754 0.754 0.870 0.746 0.745 0.760 0.766 0.760 0.806 0.782
CRPS (Norm.) 0.548 0.545 0.562 0.573 0.665 0.551 0.552 0.620 0.582 0.580 0.623 0.608
MASE (Rank) 1.333 1.667 5.667 5.667 12 3.333 3.667 7 9 7.667 11 10
CRPS (Rank) 1.667 1.333 5 6 12 3.333 3.667 10.333 7.667 7.333 10.667 9
SG PM 2.5 MASE (Raw) 0.935 0.915 0.920 0.937 0.933 0.919 0.927 0.920 0.926 0.948 0.946 0.978
CRPS (Raw) 0.320 0.311 0.317 0.319 0.317 0.312 0.315 0.335 0.320 0.326 0.326 0.339
MASE (Norm.) 0.750 0.734 0.738 0.752 0.749 0.738 0.744 0.738 0.743 0.760 0.758 0.784
CRPS (Norm.) 0.694 0.674 0.687 0.692 0.688 0.676 0.682 0.727 0.693 0.708 0.706 0.736
MASE (Rank) 7.667 2.667 3 8.667 6.333 2.667 6 3.333 6.333 10.333 9 12
CRPS (Rank) 6.333 1.667 4 6.667 5.333 2.667 4.333 11 6 9 9.333 11.667
NE China Wind MASE (Raw) 0.846 0.874 0.871 0.887 1.232 0.849 0.869 0.865 0.862 0.913 0.851 0.946
CRPS (Raw) 0.414 0.414 0.431 0.428 0.556 0.414 0.411 0.431 0.420 0.426 0.425 0.460
MASE (Norm.) 0.715 0.731 0.731 0.742 1.084 0.713 0.727 0.733 0.724 0.766 0.716 0.794
CRPS (Norm.) 0.618 0.618 0.643 0.639 0.839 0.620 0.614 0.643 0.626 0.635 0.636 0.689
MASE (Rank) 3.333 5.667 6 7.667 13 2 5.333 6.333 5 10 3.667 11
CRPS (Rank) 3.667 3 8.333 8 12 4 2.333 8.333 4.333 6.333 6.667 11
Australia Solar MASE (Raw) 0.606 0.599 0.594 0.620 0.560 0.584 0.604 0.625 0.650 0.581 0.666 0.663
CRPS (Raw) 0.232 0.227 0.233 0.253 0.229 0.236 0.239 0.268 0.243 0.235 0.247 0.259
MASE (Norm.) 0.854 0.844 0.837 0.873 0.789 0.823 0.851 0.881 0.916 0.818 0.939 0.934
CRPS (Norm.) 0.741 0.726 0.746 0.811 0.732 0.753 0.763 0.858 0.775 0.754 0.793 0.829
MASE (Rank) 6.333 5.333 4.333 8 1.333 3.333 5.667 8.667 10 2.333 11.333 11.333
CRPS (Rank) 3.333 2 4.333 10.333 2.333 5.333 6.667 12 7.333 5 8.667 10.667
EPF Electricity Price MASE (Raw) 1.221 1.193 1.272 1.207 1.474 1.163 1.229 1.206 1.232 1.235 1.243 1.289
@@
MASE (Norm.) 0.688 0.671 0.715 0.678 0.836 0.655 0.692 0.680 0.693 0.694 0.701 0.726
CRPS (Norm.) 0.637 0.630 0.665 0.643 0.793 0.618 0.671 0.683 0.654 0.656 0.647 0.677
@@
OpenElectricity NEM MASE (Raw) 0.582 0.563 0.618 0.609 0.825 0.684 0.583 0.638 0.743 0.691 0.814 1.060
CRPS (Raw) 0.174 0.160 0.191 0.183 0.240 0.185 0.165 0.193 0.186 0.183 0.205 0.238
@@
CRPS (Rank) 3 1.333 7.667 5.667 9.333 6.333 1.667 9.333 6.667 5.667 9.667 12
EWELD Load MASE (Raw) 0.667 0.650 0.686 0.677 0.776 0.653 0.696 0.667 0.669 0.675 0.853 0.994
CRPS (Raw) 0.226 0.214 0.276 0.228 0.382 0.208 0.209 0.329 0.280 0.245 0.294 0.528
MASE (Norm.) 0.673 0.654 0.695 0.682 0.789 0.660 0.702 0.673 0.678 0.680 0.864 1.009
CRPS (Norm.) 0.335 0.319 0.390 0.342 0.485 0.312 0.319 0.436 0.391 0.358 0.439 0.685
MASE (Rank) 4.333 2.667 6.667 6 9.667 2 9 4 4.667 6 11 12.667
@@
SG Carpark MASE (Raw) 0.579 0.537 0.625 0.588 0.644 0.634 0.804 0.597 0.635 0.648 1.021 0.923
CRPS (Raw) 0.040 0.037 0.045 0.041 0.045 0.043 0.052 0.044 0.046 0.045 0.067 0.067
MASE (Norm.) 0.520 0.484 0.566 0.527 0.591 0.568 0.678 0.532 0.569 0.569 0.936 0.818
CRPS (Norm.) 0.465 0.432 0.532 0.474 0.541 0.503 0.586 0.513 0.531 0.514 0.797 0.773
@@
Finland Traffic MASE (Raw) 0.522 0.505 0.540 0.525 0.560 0.607 0.594 0.563 0.596 0.559 0.895 0.639
CRPS (Raw) 0.167 0.159 0.188 0.163 0.175 0.180 0.181 0.188 0.200 0.166 0.286 0.197
MASE (Norm.) 0.555 0.536 0.576 0.558 0.600 0.647 0.633 0.590 0.636 0.595 0.946 0.682
CRPS (Norm.) 0.463 0.441 0.515 0.451 0.481 0.499 0.502 0.522 0.550 0.461 0.787 0.547
@@
CRPS (Rank) 3 1.333 7.333 3 5.333 7 6.667 8 10 4.333 12 10
Port Activity/D MASE (Raw) 0.660 0.659 0.663 0.660 0.666 0.661 0.665 0.675 0.662 0.662 0.667 0.663
CRPS (Raw) 0.591 0.599 0.619 0.588 0.619 0.589 0.594 0.699 0.596 0.599 0.601 0.595
@@
Port Activity/W MASE (Raw) 0.770 0.779 0.785 0.771 0.773 0.778 0.787 0.800 0.765 0.784 0.788 0.794
CRPS (Raw) 0.259 0.260 0.266 0.258 0.260 0.259 0.262 0.274 0.255 0.267 0.266 0.265
MASE (Norm.) 0.748 0.757 0.763 0.749 0.751 0.755 0.764 0.777 0.742 0.761 0.765 0.771
CRPS (Norm.) 0.524 0.527 0.539 0.522 0.527 0.524 0.530 0.554 0.517 0.541 0.539 0.537
@@
CRPS (Rank) 3 6 10 2 5 4 7 12 1 11 9 8
ECDC COVID/D MASE (Raw) 3.364 3.225 3.943 3.392 3.441 3.006 3.061 3.176 3.423 3.359 4.267 3.643
CRPS (Raw) 0.297 0.305 0.568 0.282 0.351 0.317 0.286 0.315 0.472 0.301 0.343 0.442
MASE (Norm.) 0.940 0.902 1.102 0.948 0.962 0.840 0.856 0.888 0.957 0.939 1.193 1.018
CRPS (Norm.) 0.479 0.492 0.916 0.456 0.566 0.511 0.462 0.509 0.762 0.485 0.552 0.713
MASE (Rank) 6 4 12 7 9 1 2 3 8 5 13 11
CRPS (Rank) 3 5 12 1 9 7 2 6 11 4 8 10
ECDC COVID/W MASE (Raw) 1.426 1.488 2.337 1.470 1.451 1.232 1.501 1.302 1.884 1.354 1.633 2.250
CRPS (Raw) 0.621 0.577 0.883 0.591 0.567 0.442 0.589 0.423 0.661 0.473 0.609 0.833
MASE (Norm.) 0.878 0.916 1.439 0.905 0.893 0.759 0.924 0.801 1.160 0.833 1.005 1.385
CRPS (Norm.) 0.780 0.725 1.110 0.743 0.713 0.556 0.740 0.532 0.831 0.594 0.765 1.047
MASE (Rank) 4 7 13 6 5 1 8 2 11 3 10 12
CRPS (Rank) 9 5 13 7 4 2 6 1 10 3 8 12
Global Influenza MASE (Raw) 0.524 0.583 0.544 0.560 0.666 0.639 0.564 0.671 0.626 0.632 0.570 0.580
CRPS (Raw) 0.374 0.363 0.395 0.344 0.684 0.401 0.396 0.429 0.467 0.452 0.348 0.425
MASE (Norm.) 0.445 0.495 0.461 0.475 0.565 0.542 0.479 0.569 0.531 0.536 0.483 0.492
CRPS (Norm.) 0.076 0.074 0.080 0.070 0.139 0.081 0.080 0.087 0.095 0.092 0.070 0.086
MASE (Rank) 1 7 2 3 11 10 4 12 8 9 5 6
CRPS (Rank) 4 3 5 1 12 7 6 9 11 10 2 8
Crypto MASE (Raw) 5.706 5.674 7.074 5.713 8.826 5.465 6.080 5.714 5.900 6.025 6.304 6.364
@@
MASE (Norm.) 1.005 1 1.246 1.006 1.555 0.963 1.071 1.007 1.039 1.061 1.111 1.121
CRPS (Norm.) 0.887 0.853 0.981 0.889 1.351 0.842 0.934 0.909 0.915 0.946 0.952 0.952

### v2

@@ Appendix E Full Results
Coastal T-S/H MASE (raw) 1.618 1.612 1.732 1.637 1.797 1.718 1.602 1.955 1.859 1.710 1.896 1.909
CRPS (raw) 0.029 0.029 0.033 0.030 0.033 0.031 0.029 0.034 0.035 0.031 0.033 0.038
MASE (norm.) 0.778 0.778 0.842 0.791 0.902 0.820 0.774 0.925 0.888 0.828 0.920 0.920
CRPS (norm.) 0.576 0.588 0.648 0.597 0.667 0.617 0.576 0.682 0.679 0.615 0.667 0.739
MASE (rank) 2.667 2.667 7.333 4 7.667 5.667 2.333 11 8.667 6 10 10.667
CRPS (rank) 2 3.667 7.667 4 8.333 5 1.333 10.667 9.333 5.667 9.333 11
SG Weather MASE (raw) 0.881 0.876 0.899 0.899 1.037 0.886 0.888 0.906 0.913 0.900 0.961 0.932
CRPS (raw) 0.177 0.175 0.181 0.185 0.215 0.177 0.178 0.200 0.188 0.186 0.201 0.196
MASE (norm.) 0.739 0.735 0.754 0.754 0.870 0.743 0.745 0.760 0.766 0.755 0.806 0.782
CRPS (norm.) 0.548 0.542 0.562 0.573 0.665 0.550 0.552 0.620 0.582 0.579 0.623 0.608
MASE (rank) 2 1 6 6 12 3 4 7.667 9 6.333 11 10
CRPS (rank) 2.333 1 5 6 12 3 3.667 10.333 7.667 7.333 10.667 9
SG PM 2.5 MASE (raw) 0.935 0.915 0.920 0.937 0.933 0.919 0.927 0.920 0.926 0.949 0.946 0.978
CRPS (raw) 0.320 0.311 0.317 0.319 0.317 0.312 0.315 0.335 0.320 0.327 0.326 0.339
MASE (norm.) 0.750 0.734 0.738 0.752 0.749 0.738 0.744 0.738 0.743 0.761 0.758 0.784
CRPS (norm.) 0.694 0.673 0.687 0.692 0.688 0.676 0.682 0.727 0.693 0.709 0.706 0.736
MASE (rank) 7.667 2 3.333 8.667 6.333 2.667 6 3.333 6.667 10.333 9 12
CRPS (rank) 6.333 1.333 4 6.667 5.333 3 4.333 11 6 9 9.333 11.667
NE China Wind MASE (raw) 0.846 0.864 0.871 0.887 1.232 0.848 0.869 0.865 0.862 0.912 0.851 0.946
CRPS (raw) 0.414 0.410 0.431 0.428 0.556 0.414 0.411 0.431 0.420 0.426 0.425 0.460
MASE (norm.) 0.715 0.722 0.731 0.742 1.084 0.712 0.727 0.733 0.724 0.765 0.716 0.794
CRPS (norm.) 0.618 0.613 0.643 0.639 0.839 0.620 0.614 0.643 0.626 0.635 0.636 0.689
MASE (rank) 3.333 4.333 6.333 7.667 13 2.333 5.667 6.333 5.333 10 3.667 11
CRPS (rank) 3.667 2 8.333 8 12 4 3 8.333 4.333 6.333 7 11
Australia Solar MASE (raw) 0.606 0.600 0.594 0.620 0.560 0.585 0.604 0.625 0.650 0.584 0.666 0.663
CRPS (raw) 0.232 0.231 0.233 0.253 0.229 0.236 0.239 0.268 0.243 0.245 0.247 0.259
MASE (norm.) 0.854 0.844 0.837 0.873 0.789 0.824 0.851 0.881 0.916 0.823 0.939 0.934
CRPS (norm.) 0.741 0.740 0.746 0.811 0.732 0.754 0.763 0.858 0.775 0.785 0.793 0.829
MASE (rank) 6.333 5.667 5 7.667 1.333 3.333 5 8.667 10 2.333 11.333 11.333
CRPS (rank) 2.667 3 3.333 10.333 2 4.333 6.333 12 7 7.667 8.667 10.667
EPF Electricity Price MASE (raw) 1.221 1.189 1.272 1.207 1.474 1.163 1.229 1.206 1.232 1.235 1.243 1.289
@@
MASE (norm.) 0.688 0.669 0.715 0.678 0.836 0.655 0.692 0.680 0.693 0.694 0.701 0.726
CRPS (norm.) 0.637 0.629 0.665 0.643 0.793 0.618 0.671 0.683 0.654 0.656 0.647 0.677
@@
OpenElectricity NEM MASE (raw) 0.582 0.564 0.618 0.609 0.825 0.685 0.583 0.638 0.743 0.691 0.814 1.060
CRPS (raw) 0.174 0.158 0.191 0.183 0.240 0.185 0.165 0.193 0.186 0.183 0.205 0.238
@@
CRPS (rank) 3 1 7.667 5.667 9.333 6.333 2 9.333 6.667 5.667 9.667 12
EWELD Load MASE (raw) 0.667 0.659 0.686 0.677 0.776 0.653 0.696 0.667 0.669 0.674 0.853 0.994
CRPS (raw) 0.226 0.210 0.276 0.228 0.382 0.208 0.209 0.329 0.280 0.245 0.294 0.528
MASE (norm.) 0.673 0.662 0.695 0.682 0.789 0.660 0.702 0.673 0.678 0.680 0.864 1.009
CRPS (norm.) 0.335 0.316 0.390 0.342 0.485 0.312 0.319 0.436 0.391 0.358 0.439 0.685
MASE (rank) 4.333 3 6.667 6 9.667 2 9 4 4.667 5.667 11 12.667
@@
SG Carpark MASE (raw) 0.579 0.529 0.625 0.588 0.644 0.641 0.804 0.597 0.635 0.648 1.021 0.923
CRPS (raw) 0.040 0.036 0.045 0.041 0.045 0.043 0.052 0.044 0.046 0.045 0.067 0.067
MASE (norm.) 0.520 0.477 0.566 0.527 0.591 0.575 0.678 0.532 0.569 0.569 0.936 0.818
CRPS (norm.) 0.465 0.422 0.532 0.474 0.541 0.507 0.586 0.513 0.531 0.514 0.797 0.773
@@
Finland Traffic MASE (raw) 0.522 0.510 0.540 0.525 0.560 0.608 0.594 0.563 0.596 0.559 0.895 0.639
CRPS (raw) 0.167 0.163 0.188 0.163 0.175 0.180 0.181 0.188 0.200 0.166 0.286 0.197
MASE (norm.) 0.555 0.542 0.576 0.558 0.600 0.647 0.633 0.590 0.636 0.595 0.946 0.682
CRPS (norm.) 0.463 0.452 0.515 0.451 0.481 0.499 0.502 0.522 0.550 0.461 0.787 0.547
@@
CRPS (rank) 3 1.667 7.333 2.667 5.333 7 6.667 8 10 4.333 12 10
Port Activity/D MASE (raw) 0.660 0.659 0.663 0.660 0.666 0.660 0.665 0.675 0.662 0.662 0.667 0.663
CRPS (raw) 0.591 0.599 0.619 0.588 0.619 0.589 0.594 0.699 0.596 0.600 0.601 0.595
@@
Port Activity/W MASE (raw) 0.770 0.779 0.785 0.771 0.773 0.775 0.787 0.800 0.765 0.783 0.788 0.794
CRPS (raw) 0.259 0.260 0.266 0.258 0.260 0.261 0.262 0.274 0.255 0.267 0.266 0.265
MASE (norm.) 0.748 0.757 0.763 0.749 0.751 0.753 0.764 0.777 0.742 0.761 0.765 0.771
CRPS (norm.) 0.524 0.527 0.539 0.522 0.527 0.529 0.530 0.554 0.517 0.541 0.539 0.537
@@
CRPS (rank) 3 5 10 2 4 6 7 12 1 11 9 8
ECDC COVID/D MASE (raw) 3.364 3.229 3.943 3.392 3.441 2.916 3.061 3.176 3.423 3.499 4.267 3.643
CRPS (raw) 0.297 0.308 0.568 0.282 0.351 0.298 0.286 0.315 0.472 0.310 0.343 0.442
MASE (norm.) 0.940 0.902 1.102 0.948 0.962 0.815 0.856 0.888 0.957 0.978 1.193 1.018
CRPS (norm.) 0.479 0.496 0.916 0.456 0.566 0.481 0.462 0.509 0.762 0.501 0.552 0.713
MASE (rank) 5 4 12 6 8 1 2 3 7 9 13 11
CRPS (rank) 3 5 12 1 9 4 2 7 11 6 8 10
ECDC COVID/W MASE (raw) 1.426 1.441 2.337 1.470 1.451 1.312 1.501 1.302 1.884 1.340 1.633 2.250
CRPS (raw) 0.621 0.547 0.883 0.591 0.567 0.541 0.589 0.423 0.661 0.493 0.609 0.833
MASE (norm.) 0.878 0.887 1.439 0.905 0.893 0.808 0.924 0.801 1.160 0.825 1.005 1.385
CRPS (norm.) 0.780 0.688 1.110 0.743 0.713 0.680 0.740 0.532 0.831 0.620 0.765 1.047
MASE (rank) 4 5 13 7 6 2 8 1 11 3 10 12
CRPS (rank) 9 4 13 7 5 3 6 1 10 2 8 12
Global Influenza MASE (raw) 0.524 0.570 0.544 0.560 0.666 0.595 0.564 0.671 0.626 0.619 0.570 0.580
CRPS (raw) 0.374 0.364 0.395 0.344 0.684 0.390 0.396 0.429 0.467 0.451 0.348 0.425
MASE (norm.) 0.445 0.484 0.461 0.475 0.565 0.505 0.479 0.569 0.531 0.525 0.483 0.492
CRPS (norm.) 0.076 0.074 0.080 0.070 0.139 0.079 0.080 0.087 0.095 0.091 0.070 0.086
MASE (rank) 1 6 2 3 11 8 4 12 10 9 5 7
CRPS (rank) 4 3 6 1 12 5 7 9 11 10 2 8
Crypto MASE (raw) 5.706 5.660 7.074 5.713 8.826 5.474 6.080 5.714 5.900 6.054 6.304 6.364
@@
MASE (norm.) 1.005 0.997 1.246 1.006 1.555 0.964 1.071 1.007 1.039 1.067 1.111 1.121
CRPS (norm.) 0.887 0.852 0.981 0.889 1.351 0.842 0.934 0.909 0.915 0.946 0.952 0.952